# https://github.com/NVlabs/OmniDrive

### v1

@@ 3 OmniDrive-nuScenes
3D Distance. Based on the randomly generated 3D coordinate, identify the traffic elements near the corresponding location and provide the 3D properties of the traffic elements.
Lane-to-objects. Based on the randomly selected lane center-line, list the objects present on this lane and their 3D properties.
3.3 Metrics
The proposed OmniDrive dataset involves captioning, open-loop planning and counterfactual reasoning tasks. Each of these tasks has distinct emphasis, and it’s challenging to evaluate them with a single metric. In this section, we will elaborate on how we assess the performance of models on our dataset.
For caption-related tasks, such as scene description and the selection of attention objects, we utilize the commonly employed language-based metrics to evaluate the sentence similarity at the word level, including METEOR [3], ROUGE [27], and CIDEr [49]. Following BEV-Planner [26], Collision Rate and Intersection Rate with the road boundary are adopted to evaluate the performance of the Open-loop planning. To evaluate the performance of the counterfactual reasoning, we ask GPT-3.5 to extract keywords based on the predictions. The keywords include ‘safety,’ ‘collision,’, ‘running a red light,’ and ‘out of the drivable area.’ Then we compare extracted keywords with the ground truth to calculate the Precision and Recall for each category of the accident.
@@ 4 Experiment
4.1 Implementation Details
During the 2D pre-training stage, the training data and strategies, including batchsize, learning rate, and optimizer are the same as LLaVA v1.5’s [30]. In the finetuning stage, the model is trained by AdamW [34] optimizer with a batch size of 16. The learning rate for the projector is 4e-4, while the visual encoder and the LLM’s learning rates are 5e-4. The cosine annealing policy is employed for the training stability. The models in the ablation study are trained for 6 epochs unless specified otherwise. The number of object queries, lane queries and carrier queries is set to 900, 300 and 256 respectively.
We also explore alternative architectures. Q-Former2D is initialized with 2D pre-trained weights. It processes the image features individually in the projector and fuses them in the LLM. The Dense BEV approach uses LSS method [41, 40] to transform perspective features into a BEV feature map. We implement temporal modeling following SOLOFusion [40]. The BEV features will be consecutively fed into a MLP projector and a LLM.
Counterfactual Open-loop
Ablation Exp. Safe Red Light Collision Drivable Area Col(%) Inter(%)
P R P R P R P R Avg. Avg.
Full Model Q-Former3D $70.7$ 49.0 57.6 $58.3$ 32.3 72.6 48.5 58.6 3.79 $4.59$
Data No Online 69.4 39.4 36.2 65.6 29.7 69.4 48.0 57.8 4.93 4.02
Architecture Q-Former2D 71.4 $39.3$ 58.3 $61.1$ $32.0$ $66.7$ $44.4$ $52.8$ $3.98$ $6.03$
Dense BEV $70.2$ $17.3$ $48.7$ $53.6$ $31.1$ 70.4 $32.4$ 56.6 $4.43$ $8.56$
No Temporal 67.8 48.4 47.0 62.6 31.2 63.8 46.5 55.3 $6.07$ $5.83$
Perception No Lane 67.7 57.3 58.1 59.6 31.0 56.7 47.9 56.8 $4.65$ $8.71$
Supervision No Object & Lane $69.0$ 57.8 $51.3$ $61.2$ $30.0$ $53.2$ $45.3$ $57.1$ $6.77$ $8.43$
Table 2: Ablation study on planning related tasks. P and R represent Precision and Recall respectively. "No online" means removing the online traing data. "No temporal" means removing the temporal modules. Freeze means no gradients are applied to the backbone layers. "No Object" and "No Lane" indicate no corresponding 3D perception supervision. In this table, none of the models in the planner are using high-level command and ego status.
4.2 Planning with Counterfactual Reasoning
Based on our Omnidrive-nuScenes, we ablate various modifications in training recipes and model architectures. All analysis shown in Tab. 2 are conducted without using high-level commands and ego status [20, 26]. Under this setting, it can be observed that there is a certain correlation between counterfactual metrics and open-loop planning.
We found that Q-former2D performs better on 2D-related tasks, such as determining the status of traffic lights. However, Q-Former3D clearly has a greater advantage in 3D tasks such as collision detection (32.2% in Precision and 72.6% in Recall) and drivable area identification (48.5% in Precision and 58.6% in Recall). The models with center-line construction tasks (i.e., Full Model) outperforms the one without lane supervision in drivable area tasks.
4.3 Ablation Study & Analysis
Counterfactual Reasoning and Captioning. In Tab. 3, the Full Model achieves the best performance in terms of the counterfactual reasoning, with average precision of 52.3% and average recall of 59.6% .
More importantly, we show that our model benefits strongly from Q-Former3D and achieves comparable performance to that of Q-Former2D on caption tasks with $38.0\%$ METEOR score, $68.6\%$ CIDEr score and $32.6$ ROUGE score. Furthermore, our model can process multi-view cameras simultaneously, while Q-Former2D processes each view separately and necessitating an inefficiently high number of tokens (1500+) for input to LLM. We note that the dense BEV model yields the poorest results because it fails to leverage the benefits brought by 2D pre-training. We also found that, with the same Q-Former equipped with pre-training, the performance on descriptive tasks is similar. Introducing additional 3D supervision and temporal information did not result in significant improvements.
Abaltion Exp. Counterfactual Caption
[…]
Full Model Q-Former3D 52.3 59.6 38.0 68.6 32.6
Data No Online 45.8 58.1 38.2 69.0 32.7
Architecture Q-Former2D 51.5 55.0 38.3 67.1 32.5
Dense BEV 45.6 49.5 35.6 59.5 27.8
No Temporal 48.1 57.5 37.9 68.4 32.6
Perception No Lane 51.2 57.6 38.0 67.8 32.6
Supervision No Object & Lane 48.9 57.3 38.2 67.8 32.6
Table 3: Analysis on our benchmark.
Table 4: Results on NuScenes-QA [43]. L and C represent Lidar and Camera respectively. We highlight the SoTA methods in each modality.
Model Modality Acc.(%)
BEVDet+BUTD [43] C 57.0
BEVDet+MCAN [43] C 57.9
CenterPoint+BUTD [43] L 58.1
CenterPoint+MCAN [43] L 59.5
OmniDrive C 59.2
Comparison on NuScenes-QA. We also present results on NuScenes-QA in Tab. 4. In NuScenes-QA, most answers in NuScenes-QA are single-word and related to perception only. In the same camera modality, our model surpasses BEVDet+MCAN by 1.3% in accuracy, demonstrating the importance of pre-training. Our model’s performance is comparable to the Lidar modality’s models.
@@ 4.3 Ablation Study & Analysis
OmniDrive - - 0.40 0.80 1.32 0.84 0.04 0.46 2.32 0.94 0.93 3.65 8.28 4.29
OmniDrive++ 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30 0.56 2.48 5.96 3.00
@@ 4.4 Discussion on Open-loop Planning
We compare the proposed OmniDrive with previous state-of-the-art vision-based planners in Tab. 5. The MLLM based open-loop planning can also achieve comparable performance to SoTA methods. However, as mentioned in BEV-Planner [26], it is observed that encoding the ego status significantly improves the metrics across all methods. Additionally, we found that the high-level command also drastically reduces the collision rate and the intersection rate. Previous methods provided high-level commands based on the relative position of the ground-truth trajectory, which imposed significant constraints on the network’s regression, thereby reducing the likelihood of deviating too far from the ground truth trajectory. We believe this design is also unreasonable, so in other experiments, we eliminated this setting.
Scene Description & Counterfactual Reasoning
User Can you summarize what the panoramic images show?
Ours The panoramic images depict a rainy day in an urban setting with wet roads reflecting the overcast sky. On the left-front, we see a parking lot entrance and a traffic light showing green, indicating the flow of traffic is permitted. The front view reveals a multi-lane road with vehicles, including a white car turning right, and traffic lights at an intersection. The right-front shows a pedestrian crossing with a traffic light, and a person waiting to cross. The left-rear view captures a traffic light at red, a pedestrian crossing, and a black car waiting. The rear view shows a large building under construction, with a crane towering above and a few vehicles on the road. The right-rear view includes a traffic light at red, a pedestrian crossing, and a road leading to a roundabout, with a black car approaching the intersection. The overall scene is one of a bustling city despite the damp weather.
User What would be the outcome if you decided to move slowly and make a right turn?
Ours If I were to move slowly and make a right turn, following the trajectory [PT, (+2.44, -0.09), (+4.91, -0.36), (+7.41, -0.84), (+9.91, -1.53), (+12.37, -2.44), (+14.76, -3.55)], it would result in leaving the drivable area, which is a dangerous and incorrect action. The trajectory indicates a turn that is too sharp and not aligned with the road layout, leading to potential collisions or driving off the road. In this scenario, the correct behavior would be to maintain my lane and follow the road’s curvature, ensuring I stay within the bounds of the drivable area and adhere to traffic rules.
Table 6: Visualization results on scene description and counterfactual reasoning.
4.5 Visualization
A challenging scene is shown in Tab. 6. OmniDrive-Agent shows impressive results on scene description and counterfactual reasoning. The model has a basic understanding of relative positions and can make counterfactual reasoning correctly based on given trajectories and safety considerations.
@@ 5.1 End-to-End Autonomous Driving
The objective of end-to-end autonomous driving is to create a fully differentiable system that spans from sensor input to control signals [61, 57, 42]. This system allows for the joint optimization of the entire system, thereby mitigating the accumulation of errors. The current technical road-map is primarily divided into two paths: open-loop autonomous driving and closed-loop autonomous driving.
In the open-loop autonomous driving, the training and evaluation processes are generally conducted on log-replayed real world datasets [43]. This approach overlooks the impact of interactions between the ego vehicle and other traffic participants, leading to cumulative errors. Pioneering work UniAD [17] and VAD [20] integrate modularized design of perception tasks such as object detection, tracking, and semantic segmentation into a unified planning framework. However, Ego-MLP [60] and BEV-Planner [26] highlight the limitations of open-loop end-to-end driving benchmarks. In these benchmarks, models may overfit the ego-status information to achieve unreasonably high performance. Researchers are addressing the challenges in open-loop evaluation by introducing closed-loop benchmarks. Recent works, e.g., MILE [15], ThinkTwice [19], VADv2 [7] leverage CARLA [11] as the simulator, which enables the creation of virtual environments with feedback from other agents. Researchers urgently need a reasonable way to evaluate end-to-end autonomous driving systems in the real world. MLLM models bridge the gap between data-driven decision-making and the user. This enables us to perform interpretable analysis and conduct counterfactual reasoning based on a specific trajectory, thereby enhancing the safety redundancy of the autonomous driving system.
5.2 Multimodal Language Models (MLLMs)
Muiltimodal language models leverage LLMs and various modalities’ encoders to successfully bridge the gap between language and other modalities and perform well on multimodal tasks ranging from visual question answer, captioning, and open-world detection. Some MLLMs such as CLIP [44] and ALIGN [18] utilize contrastive learning to create a similar embedding space for both language and vision. More recently, others such as BLIP-2 [24] explicitly targets multimodal tasks and takes multimodal inputs. For these models, there are two common techniques in order to align language and other input modalities: self-attention and cross-attention. LLaVa [31], PaLM-E [12], PaLI [8], and RT2 [62] utilize self-attention for alignment by interleaving or concatenating image and text tokens in fixed sequence lengths. However, self-attention based MLLMs are unable to handle high resolution inputs and are unsuitable for autonomous driving with multi-camera high solution images. Conversely, Flamingo [1], Qwen-VL [2], BLIP-2 [24], utilize cross-attention and are able to extract a fixed number of visual tokens regardless of image resolution. Because of this, our model utilizes Qformer architecture from BLIP-2 to handle our high resolution images.
@@ 5.3 Drive LLM-Agents and Benchmarks
Drive LLM-Agents. Given LLM/MLLMs’ high performance and ability to align modalities with language, there is a rush to incorporate MLLMs/LLMs with autonomous driving (AD). Most AD MLLMs methods attempt to create explainable autonomous driving with end-to-end learning. DriveGPT4 leverages LLMs to generate reasons for car actions while also predicting car’s next control signals [58]. Similarly, Drive Anywhere proposes a patch-aligned feature extraction for MLLMs that allow it to provide text query-able driving decisions [51]. Other works leverage MLLMs through graph-based VQA (DriveLM) [46] or chain-of-thought (CoT) design [47, 54]. They explicitly solve multiple driving tasks alongside typical MLLM tasks, such as generating scene description and analysis, prediction, and planning.
Benchmarks. To evaluate AD perception and planning, there are various datasets that capture perception, planning, steering, motion data (ONCE [37], NuPlan [5], nuScenes [4], CARLA [11], Waymo [13]). However, datasets with more comprehensive lanugage annotations are required to evaluate Drive LLM methods. Datasets focused on perception and tracking include reasoning, or descriptive like captions range from nuScenes-QA [43], NuPrompt,  [56]. HAD and Talk2Car both contain human like advice to best navigate the car [22, 9], while LaMPilot contains labels meant to evaluate transition from human commands to drive action [35]. Beyond scene descriptions, DRAMA [36] and Rank2Tell [45] focus on risk object localization. Contrastly, BDD-X, Reason2Drive focus on car explainability by providing reasons behind ego car’s action and behavior [23, 38, 39]. LingoQA [38] has introduced counterfactual questions into the autonomous driving QA dataset. We believe that the interpretability and safety redundancy of autonomous driving in the open-loop setting can be further enhanced by applying counterfactual reasoning to 3D trajectory analysis.

### v2

@@ 3.1 Omni-Q
$(Q,K,V)=(\textbf{[}Q_{c},{Q}_{d}\textbf{]},P_{m}+F_{m},F_{m}),$ (2)
$\tilde{Q}=\text{Multi-head Attention}(Q,K,V).$
After that, the perception queries $Q_{d}$ are used to predict the categories and coordinates of the foreground elements. The carrier queries $Q_{c}$ are sent to a MLP to align with the dimension of LLM tokens (e.g. 4096 dimensions in LLaMA2-7B [43]) and further used for text generation following LLaVA [25].
In Omni-Q, the carrier queries play the role of the visual-language alignment. Additionally, this design enables carrier queries to leverage the geometric priors provided by the 3D position encoding, while also allowing them to leverage query-based representations acquired through the 3D perception tasks.
3.2 Omni-L
Omni-L follows the design of LLaVA [25], utilizing a simple MLP for aligning the visual-language embedding space. We extend LLaVA’s single image input to multiple images, flattening the multi-view image features $F_{m}$ and feeding them into the large language model. To distinguish different viewpoints, we add 3D position encoding $P_{m}$ to each image patch. However, for training stability, these position encoding weights are initialized to zero.
3.3 Training strategy
The training of Omni-L/Q comprises two stages: 2D-Pretraining and 3D-Finetuning. In the initial stage, we pretrain the VLMs on 2D image tasks to initialize the Q-Former/MLP Projector. Following this, the model is fine-tuned on 3D-related driving tasks (e.g. motion planning, coutner-factual reasoning, etc.). In both stages, we calculate the text generation loss without considering contrasting learning and matching loss for in BLIP-2 [20].
@@ 4 Experiment
Omni-L+$\ddagger$ - 0.31 0.62 1.06 0.66 0.35 2.41 0.92 0.64 2.78 2.48 5.62 3.63
Omni-L - - 1.43 2.34 3.24 2.34 0.23 1.47 4.00 1.90 0.90 2.82 6.16 3.29
Omni-L++ - 0.15 0.36 0.70 0.40 0.06 0.27 0.72 0.35 0.49 1.99 4.86 2.45
@@
4.1 Implementation details
During the 2D pre-training stage, the training data and strategies, including batchsize, learning rate, and optimizer are the same as LLaVA v1.5’s [24]. In the finetuning stage, the model is trained by AdamW [29] optimizer with a batch size of 16. The learning rate for the projector is 4e-4, while the visual encoder and the LLM’s learning rates are 2e-5. The cosine annealing policy is used for training stability.
We also explore alternative architectures. The BEV-MLP approach uses LSS method [36, 35] to transform perspective features into a BEV feature map. We implement temporal modeling following SOLOFusion [35]. The BEV features will be consecutively fed into a MLP projector and a LLM. In the following section, both our models – Omni-Q and Omni-L – are trained on OmniDrive unless stated otherwise.
4.2 Dataset & metrics
OmniDrive The proposed OmniDrive dataset involves captioning, open-loop planning and counterfactual reasoning tasks. In this section, we elaborate on how we assess the performance of models on our dataset. For caption-related tasks, such as scene description and the selection of attention objects, we utilize the commonly employed language-based metrics to evaluate the sentence similarity CIDEr [44]. Following BEV-Planner [22], Collision Rate and Intersection Rate with the road boundary are adopted to evaluate the performance of Open-loop planning. To evaluate the performance of the counterfactual reasoning, we ask GPT-3.5 to extract keywords based on the predictions. The keywords include ‘safety,’ ‘collision,’, ‘running a red light,’ and ‘out of the drivable area.’ Then we compare extracted keywords with the ground truth to calculate the Precision and Recall for each category of the accident.
DriveLM The DriveLM [41] dataset is designed for end-to-end autonomous driving, featuring Graph Visual Question Answering (GVQA) to handle complex dependencies. It includes 696 scenes from the nuScenes dataset, with 4,072 samples and around 0.3 million image-question pairs. These questions cover perception, prediction, planning, and behavior, helping to fine-tune models and improve performance. DriveLM’s evaluation metrics include language metrics like BLEU, ROUGE_L, and CIDEr for text generation, accuracy for multiple-choice questions, and the ChatGPT Score for open-ended Q&A. The Match Score assesses the alignment of predicted 2D boxes with ground truth objects. The final score is the weighted combination of GPT Score (0.4), Language Score (0.2), Match Score (0.2), and Accuracy (0.2).
Dataset Acc. CG. Blue RL. Cl. Mat. Score
DriveLM 0.60 0.65 0.50 0.71 0.07 0.36 0.53
+OmniDrive 0.70 0.65 0.52 0.73 0.13 0.37 0.56
+LLaVA665k 0.76 0.63 0.53 0.73 0.15 0.37 0.57
+Both 0.78 0.64 0.54 0.73 0.15 0.37 0.58
Table 3: The performance of Omni-L on the DriveLM benchmark. We added pre-training with OmniDrive and LLaVA665k, which significantly improves performance.
Ablation Exp. Safe Red Light Collision Drivable Area
P R P R P R P R
Architecture Omni-L 72.1 58.0 59.2 63.3 34.3 71.3 49.1 59.2
Omni-Q 70.7 49.0 57.6 58.3 32.3 72.6 48.5 58.6
BEV-MLP 70.2 17.3 48.7 53.6 31.1 70.4 32.4 56.6
Perception No Lane 67.7 57.3 58.1 59.6 31.0 56.7 47.9 56.8
Supervision No Object & Lane 69.0 57.8 51.3 61.2 30.0 53.2 45.3 57.1
Table 4: Analysis on OmniDrive counterfactual reasoning and open-loop planning (without ego status). P and R represent Precision and Recall respectively. “No Object” and “No Lane” indicate no corresponding 3D perception supervision in Omni-Q.
@@ 4.3 Discussion on open-loop planning
We compare Omni-L and Omni-Q with previous SoTA vision-based planners on nuScenes open-loop planning in Tab. 2. The VLM-based open-loop planning can also achieve comparable performance to SoTA methods when using ego status. However, as mentioned in BEV-planner [22], encoding the ego status significantly improves the metrics across all methods. We found that large language models tend to overfit more easily to ego status. Omni-Q exhibits weaker VLM capabilities, making the overfitting more pronounced compared to Omni-L, as evidenced by a lower L2 error but a higher collision rate. Omni-L performs significantly better than Omni-Q without using ego status. Additionally, if the model is trained solely on trajectory prediction tasks, the distribution modeling capability of the language model can degrade, leading to poor open-loop planning results. However, training with our Q&A data without ego status effectively mitigates this issue, reducing the collision rate from 3.22% to 1.90% and the intersection rate from 3.90% to 3.29%. Critically, training Omni-L with both OmniDrive and ego status leads to significant improvements across metrics.
4.4 Results on DriveLM dataset
Our dataset, OmniDrive, plays a crucial role in pre-training. Training with only DriveLM dataset leads to an average score of 53%. However, incorporating our pre-training data boosts performance by 3%. Additionally, when combined with LLaVA665K pre-training, our data still provides a significant improvement. Although our annotations are highly automated, the OmniDrive dataset maintains high quality through counterfactual-based checklist and human-in-the-loop validation.
4.5 Planning with counterfactual reasoning
We evaluated our OmniDrive counterfactual reasoning tasks in Tab. 4. We observe that Omni-L, designed from a VLM perspective, performs better on average (e.g. 72.1% Precision and 58.0% Recall in safety tasks) compared to Omni-Q. However, Omni-Q benefits from 3D perception supervision, showing improvements in tasks like collision detection (32.3% Precision and 72.6% Recall) compared to those without 3D supervision. The combination of BEV features and MLP results in poorer performance due to pre-training gaps (e.g. 17.3% Recall in safety tasks).
Ablation Exp. Counterfactual Language Open-loop
AP (%) $\uparrow$ AR (%) $\uparrow$ CIDEr $\uparrow$ Col(%) $\downarrow$ Inter(%) $\downarrow$
Architecture Omni-L 53.7 63.0 73.2 1.90 3.29
Omni-Q 52.3 59.6 68.6 3.79 4.59
BEV-MLP 45.6 49.5 59.5 4.43 8.56
Perception No Lane 51.2 57.6 67.8 4.65 8.71
Supervision No Object & Lane 48.9 57.3 67.8 6.77 8.43
Table 5: Analysis on nuScenes open-loop planning and OmniDrive counterfactual reasoning and language ability.
4.6 Ablation study & analysis
In Tab. 5, we illustrate the performance of different architectures on OmniDrive counterfactual reasoning, language ability, and nuScenes open-loop planning. It shows a positive correlation between language ability and performance in these tasks. Omni-L excels with a counterfactual AP of 53.7%, AR of 63.0%, and a Language CIDEr score of 73.2, alongside better open-loop planning metrics (Col 1.90%, Inter 3.29%). In contrast, Omni-Q, despite benefiting from 3D perception, has lower results with a counterfactual AP of 52.3%, AR of 59.6%, and a CIDEr score of 68.6, due to weaker foundational language skills. This highlights the need for future exploration on aligning traditional 3D perception stacks with language spaces to enhance performance.
@@ 5.1 End-to-end autonomous driving
The objective of end-to-end autonomous driving is to create a fully differentiable system that spans from sensor input to control signals [55, 52, 37]. The current technical road-map is primarily divided into two paths: open-loop autonomous driving and closed-loop autonomous driving.
In the open-loop autonomous driving, the training and evaluation processes are generally conducted on log-replayed real world datasets [38]. Pioneering work UniAD [14] and VAD [17] integrate modularized design of perception tasks such as object detection, tracking, and semantic segmentation into a unified planning framework. However, Ego-MLP [54] and BEV-Planner [22] highlight the limitations of open-loop end-to-end driving benchmarks. In these benchmarks, models may overfit the ego-status information to achieve unreasonably high performance. Researchers are addressing the challenges in open-loop evaluation by introducing closed-loop benchmarks. Recent works, e.g., MILE [13], ThinkTwice [16], VADv2 [5] leverage CARLA [9] as the simulator, which enables the creation of virtual environments with feedback from other agents. Researchers urgently need a reasonable way to evaluate end-to-end autonomous driving systems in the real world. VLM models enable us to perform interpretable analysis and conduct counterfactual reasoning based on a specific trajectory, thereby enhancing the safety redundancy of the agent.
5.2 Vision-language models
vision-language models leverage LLMs and various modalities’ encoders to successfully bridge the gap between language and other modalities and perform well on multimodal tasks ranging from visual question answer, captioning, and open-world detection. Some VLMs such as CLIP [39] and ALIGN [15] utilize contrastive learning to create a similar embedding space for both language and vision. More recently, others such as BLIP-2 [20] explicitly targets multimodal tasks and takes multimodal inputs. For these models, there are two common techniques in order to align language and other input modalities: self-attention and cross-attention. LLaVa [25], PaLM-E [10], PaLI [6], and RT2 [56] utilize self-attention for alignment by interleaving or concatenating image and text tokens in fixed sequence lengths. However, self-attention based VLMs are unable to handle high resolution inputs and are unsuitable for autonomous driving with multi-camera high solution images. Conversely, Flamingo [1], Qwen-VL [2], BLIP-2 [20], utilize cross-attention and are able to extract a fixed number of visual tokens regardless of image resolution. Because of this, our model utilizes Qformer architecture from BLIP-2 to handle our high resolution images.
@@ 5.3 Drive LLM-agents and benchmarks
Drive LLM-agents. Given LLM’ high performance and ability to align modalities with language, there is a rush to incorporate VLMs/LLMs with autonomous driving (AD). Most AD VLMs methods attempt to create explainable autonomous driving with end-to-end learning. DriveGPT4 leverages LLMs to generate reasons for car actions while also predicting car’s next control signals [53]. Similarly, Drive Anywhere proposes a patch-aligned feature extraction for VLMs that allow it to provide text query-able driving decisions [47]. Other works leverage VLMs through graph-based VQA (DriveLM) [41] or chain-of-thought (CoT) design [42, 50]. They explicitly solve multiple driving tasks alongside typical VLM tasks, such as generating scene description and analysis, prediction, and planning.
Benchmarks. To evaluate AD perception and planning, there are various datasets that capture perception, planning, steering, motion data (ONCE [32], nuScenes [3], CARLA [9], Waymo [11]). However, datasets with more comprehensive lanugage annotations are required to evaluate Drive LLM methods. Datasets focused on perception and tracking include reasoning, or descriptive like captions range from nuScenes-QA [38], NuPrompt,  [51]. HAD and Talk2Car both contain human like advice to best navigate the car [19, 7], while LaMPilot contains labels meant to evaluate transition from human commands to drive action [30]. Beyond scene descriptions, DRAMA [31] and Rank2Tell [40] focus on risk object localization. Contrastly, BDD-X, Reason2Drive focus on car explainability by providing reasons behind ego car’s action and behavior [18, 33, 34]. LingoQA [33] has introduced counterfactual questions into the autonomous driving QA dataset. We believe that the interpretability and safety redundancy of autonomous driving in the open-loop setting can be further enhanced by applying counterfactual reasoning to 3D trajectory analysis.